1) Following is cost structure of an independent comic book publisher:

    Fixed cost = 75,000

    Variable cost = 8.00 per book

    Selling price = 13.00 per book

Compute the following:

  1. How many books have to be sold to make a profit of 2000 dollars? (3 points)
  2. The publisher has an estimate of demand of 5000 books. At what selling price would the publisher break even? (6 points)

#2) TransCanada Corp. is in the business of manufacturing toys using assembly lines. The time to perform each task and the tasks that must immediately precede each task are:

Task                       Precedence       Duration (seconds)

A                                 ---                                18

B                                 ---                                30

C                                 A, B                            12

D                                 B                                 6

E                                 C, D                            12

F                                  D                                 36

G                                 F                                  24

Demand per day is 560 units. The company works one 8-hour shift each day. A shift includes 2 snack breaks of 20 minutes (each). In addition, every hour there is a 6-minute rest break (i.e. 48 minutes per 8-hour shift).

  1. Compute cycle time. (3 points)
  2. For a cycle time of 45 seconds, balance the line using the Longest Operation Time first rule. (4 points)
  3. If cycle time is 45 seconds, what is the production rate per day? (2 points)

#3) A Woodwork company employed six full-time (40 hours/week) workers. In April the company had four part-time workers working 10 hours per week. In May there were 2 part-time workers and they only worked 25 hours per week. The company had revenues that averaged $60,000 per week in April and $50,000 per week in May. Assume 4 weeks in each month (April and May).

What is the percentage change in labor productivity from April to May for the company?
Answer #1

1)

Given

Fixed cost FC = 75,000

Variable cost VC = 8 per book

Selling price SP = 13 per book

a)

Profit = 2000

Profit = (SP – VC)Q – FC

Q = Quantity

2000 = (13-8)Q – 75000

2000 = 5Q – 75000

Q = 15,400

b)

Breakeven point is the point where total sales and total costs are equal

Total sales = Total costs

Total sales = Demand * Selling price

Total costs = Fixed cost + Variable cost *Demand

Given Demand = 5000

5000 * SP = 75000 + 5000*8

SP = 115000/5000

Selling price SP = 23